# thanks: Contributed equallythanks: Contributed equally

### v1

@@ IV.3 Walk Order
Shortest Hamiltonian Path. MST approach minimizes the total Hamming distance for the walks, which corresponds to the optimal walk order if control reduction is not taken into account. However, when control reduction is applied, other walk orders may be more efficient since they may allow for additional control reduction.
In general, it is difficult to calculate in advance which walk orders are better for the maximum control reduction, but from the numerical experiments we discovered that, on average, path graphs (trees with exactly two leaves) are more amenable to it than other trees. Therefore, instead of MST, one could find the Shortest Hamiltonian Path (SHP) in the same graph of the target basis states as used to build the MST.
@@
A comparison of the different walking orders presented in this section for the case of $m=n$ with control reduction is shown in Fig. 7. We empirically validate the four quantum walk order methods and find that on the prepared states, SHP requires the fewest CX gates, although it approaches the MST method as $n$ increases.
@@
Figure 7: The average number of CX gates necessary to implement a random state with given number of qubits for the considered walk orders. Each state consist of $m=n$ non-zero amplitudes and each point is the average over 1000 random states.
IV.4 Additional Optimizations
@@
IV.5 Example
Suppose we want to prepare the state
[…]
Using the algorithm described above, with sorted walk order and control reduction, one can generate a circuit shown in Fig. 8. The exact values of the coefficients $c_{1},...,c_{5}$ can be disregarded here, since they only change the angles of the rotational gates, but not the overall structure of the circuit.
Figure 8: Example circuit generated by the quantum walk state preparation algorithm presented in this work. The leading CX gates in the dashed rectangles can be removed without affecting the final state.
First, we start from the |$\ket{00001}$ state by applying an X gate to the last qubit. In the second segment of the circuit, we want to connect with the state |$\ket{00110}$. This state is different from the previous state in bits 2, 3, and 4. We arbitrarily choose qubit 2 as the target qubit and make the other bits consistent by applying CX gates to qubits 3 and 4, with control on qubit 2, which transforms the |$\ket{00110}$ state to |$\ket{00101}$. This enables the quantum walks between the states |$\ket{00001}$ and |$\ket{00101}$, which consist of the Rz gate that establishes the correct phase of $c_{1}$, and Rx gate, which establishes the correct magnitude of $c_{1}$. No controls are necessary on Rz and Rx gates here since only one state has non-zero amplitude at this point. After that, we apply the same CX gates again to transform the newly populated state back to |$\ket{00110}$.
In segment 3, we want to connect the states |$\ket{00110}$ and |$\ket{00111}$. These states are already adjacent, so no CX conjugation is necessary here. We want to make sure that Rz and Rx gates only affect |$\ket{00110}$ and not |$\ket{00001}$, so we add the minimally necessary number of controls to make it happen. In this case, it is sufficient to control on qubit 3. After these operations, the correct phase and magnitude of $c_{2}$ are established.
Continuing in the same fashion, we connect the states |$\ket{00111}$ and |$\ket{01001}$ in segment 4, and states |$\ket{01001}$ and |$\ket{01011}$ in segment 5. After this, all phases and magnitudes of the target coefficients are correct, except for the phase of $c_{5}$. To fix the phase of $c_{5}$, we choose the closest zero-amplitude state to the last populated state and interact with it via the Rz gate (no Rx gate is needed here). In this case, we chose the state |$\ket{11011}$ and control on qubits 2 and 3, where the control on qubit 2 is conjugated with X gates to implement a 0-control.
Note that the leading CX gates, marked with the dashed rectangles in Fig. 8 can be removed, since their control qubits, at the time of the gate application, are guaranteed to be in the 0-state. This optimization can only be done at the beginning of the circuit, before any CRx gates acted on the corresponding qubits, therefore we show these gates here for the purpose of demonstrating a general circuit structure.
@@ V Performance
We tested the validity of our method by preparing 1000 random quantum states for each combination of $n$ and $m$ shown in Fig. 9. Each state was prepared using the quantum walk based techniques presented in this work with control reduction and SHP walk order. Additionally, the same states were also prepared with Qiskit’s built-in prepare_state method (which is based on Ref. [31]), shown here for comparison. All circuits have been transpiled into single-qubit gates and CX basis using Qiskit’s transpiler with optimization level 3.
As can be seen from Fig. 9, Qiskit’s performance scales exponentially and is very similar regardless of the value of $m$. In fact, for $m=n^{2}$ and $m=2^{n-1}$ Qiskit produces circuits with exactly the same number of CX gates regardless of the state being prepared.
In contrast to Qiskit, our method takes advantage of the sparsity of the target state and is expected to outperform Qiskit’s built-in method for any asymptotically sparse state (i.e. $m=\mathcal{O}(poly(n))$) for sufficiently large values of $n$. However, for asymptotically dense states (i.e. $m=\mathcal{O}(2^{n})$) our method is less efficient than Qiskit for sufficiently large values of $n$.
@@
Figure 9: The average number of CX gates necessary to prepare a random state using the uniformly controlled rotation method [31] used by Qiskit’s built-in state preparation algorithm and the quantum-walk based algorithm presented in this paper for different values of $n$ and $m$.
VI Conclusions and Future Directions
In this work, we developed an algorithm that can be used to convert CTQWs on dynamic graphs that consist only of self-loops or single edges to the quantum gate model. This algorithm serves as the basis for a deterministic state preparation routine that has complexity $\mathcal{O}(nm)$ where $n$ is the number of qubits and $m$ is the number of basis states in the target state. Furthermore, we introduce multiple methods that can reduce the CX count of the state preparation algorithm and test these methods against the uniformly controlled rotation method [31] used by the Qiskit transpiler. We find that our state preparation methods require fewer CX gates when the target state has a polynomial number of non-zero amplitudes.
The quantum walks QSP framework we present posts many interesting possible avenues for future investigation. As mentioned in Sections IV.2 and IV.3, the gate count for preparing an arbitrary state can be reduced depending on the order in which the state transfer occurs. Without control reduction, the optimal sequence of walks is given by the minimum spanning tree approach. Determining the optimal sequence of basis state transfers in the presence of control reduction is more complicated and could be an interesting direction for future research. Additionally, if a target state is comprised of basis states with symmetry, it may be possible to cleverly reduce the number of gates required to create the target state by exploiting symmetry. Another interesting avenue for future work is to compare against the methods in Refs. [6, 44].
@@ VI Conclusions and Future Directions
AG developed the CTQWs to gates conversion, formulated control reduction, wrote code, gathered data, constructed proofs, and calculated complexities. RH developed the algorithm that converts CTQWs on dynamic graphs to the circuit model. CC wrote code, developed state preparation approaches, and tested circuit optimization techniques. IG developed the minimum spanning tree approach, CRz implementation of the self-loop walks, wrote code, collected data and prepared some figures. JL also developed the circuit optimization for reducing control bits. TT helped develop state preparation approaches and evaluation methodology. ZHS had the idea of creating a dictionary between single edge quantum walks and quantum circuits.

### v2

@@ IV.5 Tree Construction and Walk Order
Shortest Hamiltonian Path. The MST approach minimizes the total Hamming distance for the walks, which corresponds to the optimal walk order if control reduction is not taken into account. However, when control reduction is applied, other walk orders may be more efficient since they may allow for additional control reduction. Therefore, instead of MST, one could find the Shortest Hamiltonian Path (SHP) in the same graph of the target basis states as used to build the MST.
@@
A comparison of the different walking orders presented in this section for the case of $m=n$ is shown in Fig. 7. We empirically validate the quantum walk order methods and find that on the prepared states, the MHS methods and greedy methods require the fewest CX gates. Greedy(MHS) Combined is the greedy method on the initial ordering of MHS Linear and the better circuit between the two methods is used.
@@
IV.6 Additional Optimizations
@@ V Performance
In this section we present the comparison between the results of numerical simulations obtained for different walk orders of our single-edge state preparation method, the Merging States (MS) method introduced in [6] and Qiskit’s built-in state preparation method based on uniformly controlled rotations [5]. The error bars in the figures denote the 95% confidence interval for the estimate of the mean.
V.1 Comparison of Walk Orders
First, we compare the performance of different walk orders presented in Sec. IV.5, which is shown in Fig. 7. The greedy methods use the initial ordering specified in parenthesis. For the “Greedy Combined” results, we consider both Greedy and the method in parenthesis and choose the one with the lower CX count.
As seen from the figure, all considered walking orders have roughly similar performance, with Random, Sorted, SHP and MST being somewhat worse than others. MHS Linear and Nonlinear outperform Greedy(Sorted) on average except at 5 qubits where only MHS Linear beats greedy. As expected, Greedy(MHS) Combined performs the best on average.
Figure 7: Comparison of walk orders with $m=n$. Greedy(MHS) (Greedy(Sorted)) uses MHS Linear (Sorted) as the initial ordering. Each data point is averaged over 1000 random states.
(a) $m=n$
(b) $m=n^{2}$.
Figure 8: CX count comparison between single-edge method with the best walk order (MHS Nonlinear) presented in this paper, state-of-the-art sparse state preparation method (Merging States) and Qiskit’s built-in state preparation method for (a) $m=n$ and (b) $m=n^{2}$. Each data point is averaged over 1000 randomly generated states.
V.2 Comparison to Merging States and Qiskit
In contrast to our single-edge state preparation method, which starts from one state and gradually splits existing states to form new states, the MS method of Ref. [6] instead focuses on the inverse process of merging basis states. The MS method can be interpreted as a series of quantum walks with specified targets for the multi-control gates and optimized with the CX backward-propagation method. Note that in MS, the authors of Ref. [6] use a single multi-controlled gate, denoted CG, to implement the merging (splitting) and the phase correction in one step. However, CG can also be written as a product of CP and CRx gates.
In Figs. 8a and 8b, we compare the performance of our single-edge state preparation method with the best walk order (MHS Nonlinear) to that of the Merging States method and Qiskit’s built-in prepare_state method (which is based on Ref. [35]). Just as in the previous section, an average CX gate count for all methods is calculated for a set of 1000 randomly generated sparse states ($m=n$ and $m=n^{2}$ non-zero amplitude basis states) for every value of $n$ from 5 up to 11.
As can be seen from Figs. 8a and 8b, Qiskit’s performance scales exponentially and is very similar regardless of the value of $m$. In fact, for $m=n^{2}$ and $m=2^{n-1}$ Qiskit produces circuits with exactly the same number of CX gates regardless of the state being prepared.
In contrast to Qiskit, our method takes advantage of the sparsity of the target state and is expected to outperform Qiskit’s built-in method for any asymptotically sparse state (i.e. $m=\mathcal{O}(poly(n))$) for sufficiently large values of $n$.
Our methods outperforms not only sparsity-agnostic Qiskit’s method, but also the state-of-the-art sparse state preparation method – Merging States. As one can see from the plots, for all values of $n$, our method requires fewer CX gates.
@@ V.2 Comparison to Merging States and Qiskit
VI Conclusions and Outlook
In this work, we developed an algorithm that can be used to convert CTQWs on dynamic graphs that consist only of self-loops or single edges to the quantum gate model. This algorithm serves as the basis for a deterministic state preparation routine that has complexity $\mathcal{O}(nm)$ where $n$ is the number of qubits and $m$ is the number of basis states in the target state. Furthermore, we introduce multiple methods that can reduce the CX count of the state preparation algorithm and test these methods against the MS method of [6] and uniformly controlled rotation method [35] used by the Qiskit transpiler. We find that our state preparation methods require fewer CX gates when the target state has a polynomial number of non-zero amplitudes. In Ref. [49], the authors present another sparse state preparation method that requires only 1 ancilla. We did not directly compare to their method, but their protocol is outperformed by the MS method for low-degree polynomials $m(n)$ [49]. Since our method outperforms the MS method in these regimes, our method should outperform theirs as well.
The single-edge framework we present here posts many interesting possible avenues for future investigation. As mentioned in Sections IV.2 and IV.5, the gate count for preparing an arbitrary state can be reduced depending on the order in which the state transfer occurs. Without control reduction, the optimal sequence of walks is given by the minimum spanning tree approach. Determining the optimal sequence of basis state transfers in the presence of control reduction is more complicated and could be an interesting direction for future research. Additionally, if a target state is comprised of basis states with symmetry, it may be possible to cleverly reduce the number of gates required to create the target state by exploiting symmetry.
@@ VI Conclusions and Outlook
AG developed the CTQWs to gates conversion, formulated control reduction, wrote code, gathered data, constructed proofs, and calculated complexities. RH developed the algorithm that converts CTQWs on dynamic graphs to the circuit model. CC wrote code, developed state preparation approaches, and tested circuit optimization techniques. IG developed the minimum spanning tree approach, CRz implementation of the self-loop walks, wrote code, collected data and prepared some figures. JL also developed the circuit optimization for reducing control bits. TT helped develop state preparation approaches and evaluation methodology. ZHS had the idea of creating a dictionary between single-edge quantum walks and quantum circuits.
@@ Acknowledgments
1:Convert from CTQWs to a circuit.
2:Input:
3:$walks$: The set of walks.
4:$state$: The target state.
5:Output:
6:The circuit that implements the quantum walks.
7:
8:$walks\leftarrow$ Reverse of the walks.
9:$circ\leftarrow$ Initial empty quantum circuit.
10:for $z_{1}$, $z_{2}$, $target$, $pangle$, $xangle$ in walks do
11:   if $target$ is not specified then
12:      $target\leftarrow$The qubit index that results in the smallest number of controls.}
13:   end if
114:   {Apply the CX gates required to bring $z_{1}$ and $z_{2}$ to a Hamming distance of 1.}
15:   Determine $mhs$ for $z_{1},z_{2},target$.
16:   if $z_{2}$ is a leaf node then
17:      {Apply controlled U Gate to $circ$ (corresponds to the phase, Rx, phase gate pattern described in Section IV.4.)}
18:   else
19:      Apply controlled SU(2) Gate to $circ$.
20:   end if
21:   Update the walks with the CX gates.
22:end for
23:return inverse of $circ$.
Algorithm 1 Conversion
1:Input:
2:$basis$: A set of basis states to walk through.
3:Output:
4:A sequence of single-edge quantum walks [[$z_{1}$, $z_{2}$, $target$]i] through the provided basis states that minimizes the number of controls needed to implement the corresponding multi-controlled operations.
5:
6:$path\leftarrow$ empty.
7:$mutable\_basis\leftarrow$ copy $basis$.
18:while size of $basis>1$ do
9:   $all\_mhs\_z_{1}\leftarrow$MHS for every element in $mutable\_basis$.}
10:   $all\_diffs\leftarrow$for every element $z_{i}$ in $mutable\_basis$ get the subset $k_{i}=mutable\_basis/z_{i}$ and construct a list of list of the bits that differ with $z_{i}$.}
11:   $all\_targets\leftarrow$For every element $mhs$ in $all\_mhs\_z_{1}$ get the smallest frequency index in the corresponding $all\_diffs[mhs]$}.
12:   $all\_basis\_z_{2}\leftarrow$A list of list. For every element $z_{i}$ in $mutable\_basis$ get the subset $k_{i}\subset mutable\_basis/z_{i}$ such that $k_{i}$ is only covered by the hitting set $all\_mhs\_z_{1}[z_{i}]$ with the corresponding target $all\_targets[z_{i}]$.}
13:   $all\_mhs\_z_{2}\leftarrow$A list of list of list. For every element $basis_{i}$ in $all\_basis\_z_{2}$ get the MHS for each element in $basis_{i}$.}
14:   {Append to $path$ $[z_{1},z_{2},target]$ that corresponds to the smallest size($all\_mhs\_z_{1}[z_{1}]$)$+$size($all\_mhs\_z_{2}[z_{1}][z_{2}]$). Note that $z_{1}$ and $z_{2}$ are in terms of the original basis.}
15:   Remove $z_{2}$ from $basis$.
16:   Remove $z_{2}$ from $mutable\_basis$.
217:   Update $mutable\_basis$ with CX gates.
18:end while
19:return the inverse of $path$.
Algorithm 2 MHS Nonlinear
1:The walks considered here all form linear graphs. Greedy finds the best position to insert the next basis state into the path sequence.
2:Input:
3:$basis$: Basis in some order.
4:Output:
5:The best circuit found.
16:
7:$path\leftarrow$ Empty list
8:for $z$ in $basis$ do
9:   $cxcount\leftarrow$ None.
10:   $curr\_path\leftarrow$ None.
11:   for $idx$ in (length($path$)+1) do
12:      $temp\_path\leftarrow$ copy $path$.
13:      Insert $z$ in $temp\_path$ at $idx$.
14:      $temp\_cxcount\leftarrow$ CX count for constructed circuit with $temp\_path$.
15:      if $cxcount$ is None or $temp\_cxcount<cxcount$ then
16:         $curr\_path\leftarrow temp\_path$.
17:      end if
18:   end for
19:   $path\leftarrow curr\_path$.
20:end for
21:return circuit for $path$.
Algorithm 3 Greedy Insertion